# Anytime-valid, Bayes-assisted, Prediction-Powered Inference

### v1

@@ 3.2 Asymptotic Bayes-assisted confidence sequences
To address this, we introduce a Bayes-assisted analogue of Theorem 3.3.
Theorem 3.5 (Bayes-assisted AsympCS - iid).
@@
Let $(Y_{t})_{t\geq 1}$ be a sequence of i.i.d. random variables with mean $\mu$ and finite variance. For $t\geq 1$, let $\overline{Y}_{t}$ be the sample mean, and $\widehat{\sigma}_{t}^{2}$ be the sample variance based on the first $t$ observations. Let $\eta_{t}:\mathbb{R}\to(0,\sqrt{t/(2\pi)})$ be defined as
[…]
@@
where $\pi$ be a continuous and proper prior density on $\mathbb{R}$. Then
@@
$\mathcal{C}^{\mathrm{\mathtt{B}A}}_{\alpha,t}(\overline{Y}_{t},\widehat{\sigma}_{t};\pi):=\left[\overline{Y}_{t}\pm\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}\eta_{t}(\overline{Y}_{t}/\widehat{\sigma}_{t})^{2}}\right)}\right]$ (8)
forms a $(1-\alpha)$–AsympCS for $\mu$.
@@
In Theorem 3.5, the density $\pi$ encodes prior beliefs about the ratio $\mu/\sigma$. Under this prior, $\eta_{t}$ represents the marginal density of the standardized mean $\overline{Y}_{t}/\sigma$ that would arise if the observations $(Y_{t})_{t\geq 1}$ were normally distributed. In contrast to the non-Bayesian AsympCS (5), the width of the Bayes-assisted AsympCS (8) varies with $\overline{Y}_{t}/\widehat{\sigma}_{t}$: when the data align with the prior, $\eta_{t}(\overline{Y}_{t}/\widehat{\sigma}_{t})$ is large and the interval narrows; when they conflict, $\eta_{t}(\overline{Y}_{t}/\widehat{\sigma}_{t})$ is small and the interval widens. It is worth emphasizing that, even when if prior is strongly misspecified, the Bayes-assisted AsympCS (8) remains valid. In the case of a Gaussian prior $\pi$ centered at $\mu_{0}$ with variance $\tau^{2}$, we obtain the following AsympCS :
@@
$\mathcal{C}^{\mathrm{\mathtt{B}A}}_{\alpha,t}(\overline{Y}_{t},\widehat{\sigma}_{t};\mathcal{N}(\cdot;\mu_{0},\tau^{2}))=\left[\overline{Y}_{t}\pm\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t\tau^{2}+1}{\alpha^{2}}\right)+\frac{(\overline{Y}_{t}/\widehat{\sigma}_{t}-\mu_{0})^{2}}{(\tau^{2}+1/t)}}\right].$ (9)
@@
Setting $\rho=\tau$ allows a direct comparison between (9) and its non-Bayesian counterpart (5). When the data agree with the prior – that is, $Y_{t}/\widehat{\sigma}_{t}-\mu_{0}\simeq 0$ – the Bayes-assisted interval is narrower than the non-assisted one. Conversely, if the data conflict with the prior, $(Y_{t}/\widehat{\sigma}_{t}-\mu_{0})^{2}$ is large and the Bayes-assisted AsympCS becomes wider than (5).
@@
The proof of Theorem 3.5 is similar to that of (Waudby-Smith et al., 2024b, Theorem 2.2). First, note that $\mathcal{C}^{\mathrm{\mathtt{B}A}}_{\alpha,t}(\overline{Y}_{t},\mathrm{var}(Y);\pi)$ would be an exact CS if the observations were normally distributed. This follows from an application of the method of mixture for nonnegative martingales, using the prior $\pi$ as mixing density, together with Ville’s inequality. Second, we use Strassen’s strong coupling to approximate in an almost sure sense $\overline{Y}_{t}$ by a sample average of i.i.d. Gaussian random variables. As in the non-Bayesian case, Theorem 3.5 can be extended to the non-iid setting, as long as one can find such strong coupling.
Theorem 3.6 (Asymptotic Bayes-assisted CS - non-iid case).
@@
Consider the same notation and assumptions as in Theorem 3.4. Let $\pi$ be a continuous and proper prior density on $\mathbb{R}$ and let $\eta_{t}$ be the density (7) for any $t\geq 1$. Then, the sequence of intervals $(\mathcal{C}^{\mathrm{\mathtt{B}A}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\sigma}_{t};\pi))_{t\geq 1}$ forms a $(1-\alpha)$–AsympCS for $\mu$.
@@ 4 Control variates and PPI: background and strong coupling
Prediction-powered inference (PPI) closely relates to control variates, a standard variance-reduction method in Monte Carlo estimation (Glasserman, 2003, §4.1). In fact, each PPI estimator can be expressed as a control-variate estimator. We begin with a short review of control variates and derive a Strassen-type strong-coupling result for these estimators, and then provide additional background on PPI.
4.1 Control variates: definitions and Strassen’s coupling
@@
where $\overline{U}$ and $\overline{V}$ denote the empirical means of $(U_{i})_{i=1}^{n}$ and $(V_{i})_{i=1}^{n}$, respectively, $\lambda\in\mathbb{R}$ is a tunable coefficient, and the term $U_{i}-\mu$ acts as a control variate. The estimator $\widehat{\gamma}^{\mathrm{icv}}_{\lambda}$ is unbiased, consistent, and has variance $\mathrm{var}(\widehat{\gamma}^{\mathrm{icv}}_{\lambda})=(\mathrm{var}(V)-2\lambda\mathrm{cov}(U,V)+\lambda^{2}\mathrm{var}(U))/n$. Compared to the standard sample mean estimator $\overline{V}$, which attains variance $\mathrm{var}(\overline{V})=\mathrm{var}(V)/n$, using $\widehat{\gamma}^{\mathrm{icv}}_{\lambda}$ results in variance reduction when $\lambda<2\mathrm{cov}(U,V)/\mathrm{var}(U)$. The minimum variance is achieved at the optimal coefficient $\lambda^{*}=\mathrm{cov}(U,V)/\mathrm{var}(U)$, for which $\mathrm{var}(\widehat{\gamma}^{\mathrm{icv}}_{\lambda^{*}})=(1-\rho_{U,V}^{2})\mathrm{var}(\overline{V})$, where $\rho_{U,V}$ is the correlation between $U$ and $V$. That is, stronger correlation leads to greater variance reduction.
@@
In practice, both $\mu$ and $\lambda^{*}$ are typically unknown. When this is the case, given an additional i.i.d. sample $(\widetilde{U}_{j})_{j=1}^{N_{n}}$, where $\widetilde{U}_{1}$ has the same distribution as $U$, one can estimate $\mu$ by $\widehat{\mu}=\frac{1}{N_{n}}\sum_{j=1}^{N_{n}}\widetilde{U}_{j}$. and plug it into Equation 10. For a fixed $\lambda$, this gives
$\widehat{\gamma}_{\lambda}^{\mathrm{cv}}=\overline{V}-\lambda(\overline{U}-\widehat{\mu})=\frac{1}{n}\sum_{i=1}^{n}\left(V_{i}-\lambda(U_{i}-\widehat{\mu})\right).$ (11)
Similarly, $\lambda^{*}$ may be estimated from data as $\widehat{\lambda}=\widehat{\mathrm{cov}}((U_{i},V_{i})_{i=1}^{n})/\widehat{\mathrm{var}}((U_{i})_{i=1}^{n})$, where $\widehat{\mathrm{var}}(\cdot)$ and $\widehat{\mathrm{cov}}(\cdot)$ denote the sample variance and covariance, respectively. Plugging $\widehat{\lambda}$ into (11) defines $\widehat{\gamma}^{\mathrm{cv+}}:=\widehat{\gamma}_{\widehat{\lambda}}^{\mathrm{cv}}$, which is similar to the semi-supervised least squares estimator of Zhang et al. (2019, Eq. (2.15)). As discussed in Section 3, deriving an AsympCS requires a strong coupling between the estimator and a sequence of i.i.d. Gaussian random variables. We now establish this coupling, a key ingredient for AsympCS for CVEs (and, in particular, for PPI estimators).
Proposition 4.1 (Asymptotics for CVEs).
We have, almost surely as $n\to\infty$,
@@
$\widehat{\gamma}^{\mathrm{cv+}}$ $=\widehat{\gamma}^{\mathrm{cv}}_{\lambda^{*}}+o\left(\sqrt{\frac{\log\log n}{n}}\right)=\overline{V}-\lambda^{*}(\overline{U}-\widehat{\mu})+o\left(\sqrt{\frac{\log\log n}{n}}\right).$ (12)
Proposition 4.2 (Strassen’s coupling for CVEs).
@@
Assume that $|\frac{n}{N_{n}}-r|=O(1/\log n)$ for some $r\in[0,1]$. There exists i.i.d. Gaussian random variables $(W^{\mathrm{cv}}_{i})_{i\geq 1}$ with mean $\gamma$ and variance $\nu_{\lambda}^{\mathrm{cv}}:=\mathrm{var}(V-\lambda U)+r\mathrm{var}(\lambda U)=\mathrm{var}(V)-2\lambda\mathrm{cov}(U,V)+\lambda^{2}(1+r)\mathrm{var}(U)$ such that, almost surely as $n\to\infty$,  $\widehat{\gamma}_{\lambda}^{\mathrm{cv}}=\frac{1}{n}\sum_{i=1}^{n}W^{\mathrm{cv}}_{i}+o\left(\sqrt{\frac{\log\log n}{n}}\right).$ Similarly, there exists i.i.d. Gaussian random variables $(W^{\mathrm{cv+}}_{i})_{i\geq 1}$ with mean $\gamma$ and variance $\nu^{\mathrm{cv+}}:=\nu_{\lambda^{*}}^{\mathrm{cv}}=\mathrm{var}(V)\left[1-(1-r)\rho_{U,V}^{2}\right]$ such that, almost surely as $n\to\infty$$\widehat{\gamma}^{\mathrm{cv+}}=\frac{1}{n}\sum_{i=1}^{n}W^{\mathrm{cv+}}_{i}+o\left(\sqrt{\frac{\log\log n}{n}}\right).$ The estimators
@@
$\widehat{\nu}_{\lambda}^{\mathrm{cv}}((U_{i},V_{i})_{i=1}^{n},(\widetilde{U}_{i})_{i=1}^{N_{n}})$ $=\frac{1}{n-2}\sum_{i=1}^{n}(V_{i}-\overline{V}-\lambda(U_{i}-\overline{U}))^{2}+\frac{n\lambda^{2}}{N_{n}(N_{n}-1)}\sum_{j=1}^{N_{n}}(\widetilde{U}_{j}-\widehat{\mu})^{2}$ (13)
@@
are consistent estimators of $\nu^{\mathrm{cv}}$ and $\nu^{\mathrm{cv+}}$, respectively.
@@ 4.2 PPI estimators: definitions and asymptotic properties
Owing to Equation 2, the PPI estimator $\widehat{\theta}_{n}$ is the value of $\theta$ that solves the equation $\widehat{g}_{\theta,n}=0$, where $\widehat{g}_{\theta,n}=\widehat{m}_{\theta,n}+\widehat{\Delta}_{\theta,n}$ is an estimator of $g_{\theta}$. Here, $\widehat{m}_{\theta,n}$ and $\widehat{\Delta}_{\theta,n}$ are estimators of $m_{\theta}$ and $\Delta_{\theta}$, respectively. A typical choice for $\widehat{m}_{\theta,n}$ is the sample mean of the unlabelled data,
$\widehat{m}_{\theta,n}$ $=\frac{1}{N_{n}}\sum_{i=1}^{N_{n}}\ell_{\theta}^{\prime}(\widetilde{X}_{i},f(\widetilde{X}_{i})).$ (15)
@@
Angelopoulos et al. (2023a) use the sample mean
@@
Combining Equation 16 with Equation 15,
$\widehat{g}^{\mathrm{\texttt{PP}}}_{\theta,n}=\widehat{m}_{\theta,n}+\widehat{\Delta}^{\mathrm{\texttt{PP}}}_{\theta,n}=\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{\theta}^{\prime}(X_{i},Y_{i})\right]-\left(\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{\theta}^{\prime}(X_{i},f(X_{i}))\right]-\widehat{m}_{\theta}\right)$ (17)
is a CVE, with control variate $\ell_{\theta}^{\prime}(X_{i},f(X_{i}))-\widehat{m}_{\theta}$ and control-variate parameter $\lambda=1$. For the squared loss, the estimator $\widehat{\theta}_{n}^{\mathrm{\texttt{PP}}}$ solving $\widehat{g}^{\mathrm{\texttt{PP}}}_{\theta,n}=0$ also takes the control-variate form
$\widehat{\theta}_{n}^{\mathrm{\texttt{PP}}}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}-\left(\frac{1}{n}\sum_{i=1}^{n}f(X_{i})-\frac{1}{N_{n}}\sum_{j=1}^{N_{n}}f(\widetilde{X}_{j})\right)$ (18)
@@
Angelopoulos et al. (2023b) extend the standard PPI estimator (17) by allowing the control-variate parameter $\lambda$, which they call power-tuning parameter, to take values other than $1$. The resulting estimator is
@@
where $\widehat{\lambda}_{\theta,n}$ is the estimator $\widehat{\lambda}_{\theta,n}=\widehat{\mathrm{cov}}\left((\ell_{\theta}^{\prime}(X_{i},Y_{i}),\ell_{\theta}^{\prime}(X_{i},f(X_{i})))_{i=1}^{n}\right)/\widehat{\mathrm{var}}\left((\ell_{\theta}^{\prime}(X_{i},f(X_{i})))_{i=1}^{n}\right)$.
In this case, $\widehat{\Delta}^{\mathrm{\texttt{PP+}}}_{\theta,n}$ is a CVE with centered control variate $\ell_{\theta}^{\prime}(X_{i},f(X_{i}))-\widehat{m}_{\theta,n}$, which depends only on the black-box predictions. As a result of this choice,
$\widehat{g}^{\mathrm{\texttt{PP+}}}_{\theta,n}=\widehat{m}_{\theta,n}+\widehat{\Delta}^{\mathrm{\texttt{PP+}}}_{\theta,n}=\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{\theta}^{\prime}(X_{i},Y_{i})\right]-\widehat{\lambda}_{\theta,n}\left(\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{\theta}^{\prime}(X_{i},f(X_{i}))\right]-\widehat{m}_{\theta}\right)$ (20)
@@
Standard asymptotic confidence intervals for PPI and PPI++ rely on CLTs for the estimators $\widehat{g}_{\theta,n}$, $\widehat{m}_{\theta,n}$ and $\widehat{\Delta}_{\theta,n}$. In contrast, constructing asymptotic confidence sequences requires almost sure approximations by averages of i.i.d. Gaussian variables. As the estimators for $g_{\theta}$, $m_{\theta}$ and $\Delta_{\theta}$ are all CVEs, the asymptotic results of Proposition 4.1 and Strassen’s coupling of Proposition 4.2 both apply.

### v2

@@ 3.2 Asymptotic Bayes-assisted confidence sequences
To address this, we introduce a Bayes-assisted analogue of Theorem˜1.
Theorem 3 (Bayes-assisted AsympCS – i.i.d. case).
@@
Let $(Y_{t})_{t\geq 1}$ be a sequence of i.i.d. random variables with unknown mean $\mu$ and unknown variance $\sigma^{2}$, and such that $\mathbb{E}|Y_{1}|^{2+\delta}<\infty$ for some $\delta>0$. For any $t\geq 1$, let $\overline{Y}_{t}$ be the sample mean, and $\widehat{\sigma}_{t}^{2}$ be the sample variance based on the first $t$ observations. Let $\eta_{t}:\mathbb{R}\to(0,\sqrt{t/(2\pi)})$ be defined as
$\eta_{t}(z)=\int_{-\infty}^{\infty}\mathcal{N}\left(z;\zeta,1/t\right)\uppi(\zeta)d\zeta.$ (8)
@@
where $\uppi$ is a continuous and proper prior density on $\mathbb{R}$, strictly positive in a neighbourhood of $\mu/\sigma$. Then
@@
$\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\overline{Y}_{t},\widehat{\sigma}_{t};\uppi):=\left[\overline{Y}_{t}\pm\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}\eta_{t}(\overline{Y}_{t}/\widehat{\sigma}_{t})^{2}}\right)}\right]$ (9)
forms a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$ for $\mu$.
@@
In Theorem˜3, the density $\uppi$ encodes prior beliefs about the ratio $\mu/\sigma$. Under this prior, $\eta_{t}$ represents the marginal density of the standardised mean $\overline{Y}_{t}/\sigma$ that would arise if the observations $(Y_{t})_{t\geq 1}$ were normally distributed. In contrast to the non-assisted AsympCS (6), the width of the Bayes-assisted AsympCS (9) varies with $\overline{Y}_{t}/\widehat{\sigma}_{t}$: when the data align with the prior, $\eta_{t}(\overline{Y}_{t}/\widehat{\sigma}_{t})$ is large and the interval narrows; when they conflict, $\eta_{t}(\overline{Y}_{t}/\widehat{\sigma}_{t})$ is small and the interval widens. It is worth emphasising that, even when the prior is strongly misspecified, the Bayes-assisted AsympCS (9) remains valid. In the case of a Gaussian prior $\uppi$ centred at $\mu_{0}$ with variance $\tau^{2}$, we obtain the following AsympCS :
@@
$\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\overline{Y}_{t},\widehat{\sigma}_{t};\mathcal{N}(\cdot;\mu_{0},\tau^{2}))=\left[\overline{Y}_{t}\pm\frac{\widehat{\sigma}_{t}}{\sqrt{t}}\sqrt{\log\left(\frac{t\tau^{2}+1}{\alpha^{2}}\right)+\frac{(\overline{Y}_{t}/\widehat{\sigma}_{t}-\mu_{0})^{2}}{\tau^{2}+1/t}}\right].$ (10)
@@
Setting $\rho=\tau$ allows a direct comparison between (10) and its non-assisted counterpart (6). When the data agree with the prior – i.e., $\overline{Y}_{t}/\widehat{\sigma}_{t}-\mu_{0}\simeq 0$ – the Bayes-assisted interval is narrower than the non-assisted one. Conversely, if the data conflict with the prior, $(\overline{Y}_{t}/\widehat{\sigma}_{t}-\mu_{0})^{2}$ is large and the Bayes-assisted AsympCS becomes wider than (6).
@@
The proof of Theorem˜3 is similar to that of [15, Theorem 2.2]. First, note that $\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\overline{Y}_{t},\mathrm{var}(Y);\uppi)$ would be an exact CS if the observations were normally distributed. This follows from an application of the method of mixtures for nonnegative martingales, using the prior $\uppi$ as mixing density, together with Ville’s inequality. Second, we use KMT strong coupling to approximate in an almost sure sense $\overline{Y}_{t}$ by a sample average of i.i.d. Gaussian random variables. As in the non-assisted case, Theorem˜3 can be extended to the non-i.i.d. setting, as long as one can find such a strong coupling.
Theorem 4 (Asymptotic Bayes-assisted CS – non-i.i.d. case).
@@
Consider the same notation and assumptions as in Theorem˜2. Let $\uppi$ be a continuous and proper prior density on $\mathbb{R}$, strictly positive in a neighbourhood of $\mu/\sigma$, and let $\eta_{t}$ be the density (8) for any $t\geq 1$. Then, the sequence of intervals $(\mathcal{C}^{\mathrm{\tt BA}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\sigma}_{t};\uppi))_{t\geq 1}$ forms a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$ for $\mu$.
3.3 Asymptotic Type-I error control
The asymptotic confidence sequences defined above satisfy an asymptotic version of time-uniform Type-I error control (in the sense of [15, §2.5]; see also [24]).
Theorem 5 (Asymptotic Type-I error control).
Assume the hypotheses of one of Theorems˜1, 2, 3 and 4, and let $(\mathcal{C}_{\alpha,t})$ be the corresponding $(1-\alpha)$–AsympCS for $\mu$. Then
$\underset{m\rightarrow\infty}{\lim\inf}\Pr\left(\mu\in\mathcal{C}_{\alpha,t}\text{ ~for all }t\geq m\right)\geq 1-\alpha.$ (11)
@@ 4 Control variates and PPI: background and strong coupling
Prediction-powered inference (PPI) closely relates to control variates, a standard variance-reduction method in Monte Carlo estimation [25, §4.1]. In fact, each PPI estimator can be expressed as a control-variate estimator. We begin with a review of control variates and derive a KMT-type strong-coupling result for these estimators, and then provide additional background on PPI.
4.1 Control variates: definitions and KMT strong coupling
@@
where $\overline{U}$ and $\overline{V}$ denote the empirical means of $(U_{i})_{i=1}^{n}$ and $(V_{i})_{i=1}^{n}$, respectively, $\lambda\in\mathbb{R}$ is a tunable coefficient, and the term $U_{i}-\mu$ acts as a control variate. The estimator $\widehat{\gamma}^{\mathrm{icv}}_{\lambda}$ is unbiased, consistent, and has variance $\mathrm{var}(\widehat{\gamma}^{\mathrm{icv}}_{\lambda})=(\mathrm{var}(V)-2\lambda\mathrm{cov}(U,V)+\lambda^{2}\mathrm{var}(U))/n$. Compared to the standard sample mean estimator $\overline{V}$, which attains variance $\mathrm{var}(\overline{V})=\mathrm{var}(V)/n$, using $\widehat{\gamma}^{\mathrm{icv}}_{\lambda}$ results in variance reduction when $\lambda<2\mathrm{cov}(U,V)/\mathrm{var}(U)$. The minimum variance is achieved at the optimal coefficient $\lambda^{\star}=\mathrm{cov}(U,V)/\mathrm{var}(U)$, for which $\mathrm{var}(\widehat{\gamma}^{\mathrm{icv}}_{\lambda^{\star}})=(1-\rho_{U,V}^{2})\mathrm{var}(\overline{V})$, where $\rho_{U,V}$ is the correlation between $U$ and $V$. That is, stronger correlation leads to greater variance reduction.
@@
In practice, both $\mu$ and $\lambda^{\star}$ are typically unknown. When this is the case, given an additional i.i.d. sample $(\widetilde{U}_{j})_{j=1}^{N_{n}}$, independent of ${(U_{i},V_{i})}_{i=1}^{n}$, where $\widetilde{U}_{1}$ has the same distribution as $U$, one can estimate $\mu$ by $\widehat{\mu}=\frac{1}{N_{n}}\sum_{j=1}^{N_{n}}\widetilde{U}_{j}$ and plug it into Equation˜12. For fixed $\lambda$, this gives
[…]
Similarly, $\lambda^{\star}$ may be estimated from data as $\widehat{\lambda}=\widehat{\mathrm{cov}}((U_{i},V_{i})_{i=1}^{n})/\widehat{\mathrm{var}}((U_{i})_{i=1}^{n})$,
[…]
Proposition 1 (Asymptotics for CVEs).
Assume $\mathbb{E}|U|^{2+\delta}$ and $\mathbb{E}|V|^{2+\delta}<\infty$ for some $0<\delta<1$. Then, almost surely as $n\to\infty$,
@@
$\widehat{\gamma}^{\mathrm{cv+}}$ $=\widehat{\gamma}^{\mathrm{cv}}_{\lambda^{\star}}+o\left(\frac{1}{\sqrt{n\log n}}\right)=\overline{V}-\lambda^{\star}(\overline{U}-\widehat{\mu})+o\left(\frac{1}{\sqrt{n\log n}}\right).$ (14)
Proposition 2 (KMT coupling for CVEs).
@@
Assume $\mathbb{E}|U|^{2+\delta}$ and $\mathbb{E}|V|^{2+\delta}<\infty$ for some $0<\delta<1$. Assume additionally that $|\frac{n}{N_{n}}-r|=O(1/n^{1-a})$ with $0<a<2/(2+\delta)$, for some $r\in[0,1]$. Then, there exist i.i.d. Gaussian random variables $(W^{\mathrm{cv}}_{i})_{i\geq 1}$ with mean $\gamma$ and variance
such that, almost surely as $n\to\infty$,
@@
$\widehat{\gamma}_{\lambda}^{\mathrm{cv}}=\frac{1}{n}\sum_{i=1}^{n}W^{\mathrm{cv}}_{i}+o\left(\frac{1}{\sqrt{n\log n}}\right).$ (15)
Similarly, there exist i.i.d. Gaussian random variables $(W^{\mathrm{cv+}}_{i})_{i\geq 1}$ with mean $\gamma$ and variance $\nu^{\mathrm{cv+}}:=\nu_{\lambda^{\star}}^{\mathrm{cv}}=\mathrm{var}(V)\left[1-(1-r)\rho_{U,V}^{2}\right]$ such that, almost surely as $n\to\infty$,
$\widehat{\gamma}^{\mathrm{cv+}}=\frac{1}{n}\sum_{i=1}^{n}W^{\mathrm{cv+}}_{i}+o\left(\frac{1}{\sqrt{n\log n}}\right).$ (16)
The estimators
$\widehat{\nu}_{\lambda}^{\mathrm{cv}}((U_{i},V_{i})_{i=1}^{n},(\widetilde{U}_{j})_{j=1}^{N_{n}})$ $=\frac{1}{n-2}\sum_{i=1}^{n}(V_{i}-\overline{V}-\lambda(U_{i}-\overline{U}))^{2}+\frac{n\lambda^{2}}{N_{n}(N_{n}-1)}\sum_{j=1}^{N_{n}}(\widetilde{U}_{j}-\widehat{\mu})^{2}$ (17)
@@
are consistent estimators of $\nu_{\lambda}^{\mathrm{cv}}$ and $\nu^{\mathrm{cv+}}$, respectively, where $\widehat{\mu}=\frac{1}{N_{n}}\sum_{j=1}^{N_{n}}\widetilde{U}_{j}$.
@@ 4.2 PPI estimators: definitions and asymptotic properties
Owing to Equation˜2, the PPI estimator $\widehat{\theta}_{n}$ is the value of $\theta$ that solves the equation $\widehat{g}_{\theta,n}=0$, where $\widehat{g}_{\theta,n}=\widehat{m}_{\theta,n}+\widehat{\Delta}_{\theta,n}$ is an estimator of $g_{\theta}$. Here, $\widehat{m}_{\theta,n}$ and $\widehat{\Delta}_{\theta,n}$ are estimators of $m_{\theta}$ and $\Delta_{\theta}$, respectively. A typical choice for $\widehat{m}_{\theta,n}$ is the sample mean of the unlabelled data,
$\widehat{m}_{\theta,n}$ $=\frac{1}{N_{n}}\sum_{j=1}^{N_{n}}\ell_{\theta}^{\prime}(\widetilde{X}_{j},f(\widetilde{X}_{j})).$ (19)
@@
Angelopoulos et al. [1] use the sample mean
@@
as an estimator for $\Delta_{\theta}$. Combining Equation˜20 with Equation˜19,
$\widehat{g}^{\mathrm{\texttt{PP}}}_{\theta,n}=\widehat{m}_{\theta,n}+\widehat{\Delta}^{\mathrm{\texttt{PP}}}_{\theta,n}=\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{\theta}^{\prime}(X_{i},Y_{i})\right]-\left(\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{\theta}^{\prime}(X_{i},f(X_{i}))\right]-\widehat{m}_{\theta,n}\right)$ (21)
is a CVE, with control variate $\ell_{\theta}^{\prime}(X_{i},f(X_{i}))-\widehat{m}_{\theta,n}$ and control-variate parameter $\lambda=1$. For the squared loss, the estimator $\widehat{\theta}_{n}^{\mathrm{\texttt{PP}}}$ solving $\widehat{g}^{\mathrm{\texttt{PP}}}_{\theta,n}=0$ also takes the control-variate form
$\widehat{\theta}_{n}^{\mathrm{\texttt{PP}}}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}-\left(\frac{1}{n}\sum_{i=1}^{n}f(X_{i})-\frac{1}{N_{n}}\sum_{j=1}^{N_{n}}f(\widetilde{X}_{j})\right),$ (22)
@@
Angelopoulos et al. [7] extend the standard PPI estimator (21) by allowing the control-variate parameter $\lambda$, which they call power-tuning parameter, to take values other than $1$. The resulting estimator is
@@
where $\widehat{\lambda}_{\theta,n}$ is the estimator $\widehat{\lambda}_{\theta,n}=\widehat{\mathrm{cov}}\left((\ell_{\theta}^{\prime}(X_{i},Y_{i}),\ell_{\theta}^{\prime}(X_{i},f(X_{i})))_{i=1}^{n}\right)/\widehat{\mathrm{var}}\left((\ell_{\theta}^{\prime}(X_{i},f(X_{i})))_{i=1}^{n}\right)$. In this case, $\widehat{\Delta}^{\mathrm{\texttt{PP+}}}_{\theta,n}$ is a CVE with centred control variate $\ell_{\theta}^{\prime}(X_{i},f(X_{i}))-\widehat{m}_{\theta,n}$, which depends only on the black-box predictions. As a result of this choice,
$\widehat{g}^{\mathrm{\texttt{PP+}}}_{\theta,n}=\widehat{m}_{\theta,n}+\widehat{\Delta}^{\mathrm{\texttt{PP+}}}_{\theta,n}=\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{\theta}^{\prime}(X_{i},Y_{i})\right]-\widehat{\lambda}_{\theta,n}\left(\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{\theta}^{\prime}(X_{i},f(X_{i}))\right]-\widehat{m}_{\theta,n}\right)$ (24)
@@
Standard asymptotic confidence intervals for PPI and PPI++ rely on CLTs for the estimators $\widehat{g}_{\theta,n}$, $\widehat{m}_{\theta,n}$ and $\widehat{\Delta}_{\theta,n}$. In contrast, constructing asymptotic confidence sequences requires almost sure approximations by averages of i.i.d. Gaussian variables. Since the estimators for $g_{\theta}$, $m_{\theta}$ and $\Delta_{\theta}$ are all CVEs, the asymptotic results of Proposition˜1 and the KMT coupling of Proposition˜2 both apply.